POYO:

The graph of f(x)=x^2 was transformed to create the graph of g(x)=8(x+5)^2. Complete the sentence. The graph of g(x) shows _____ by a factor of _______ _______ by _____ units.

POYO:

ScamTheMan:

Would you be able to post the answer choices below?

Angle:

This problem uses the idea of a ` vertex form ` The vertex form of a parabola is as follows: ` y = a(x - h)^2 + k ` where ` a ` is how much it stretches or shrinks ` h ` is how much right or left it moves/shifts ` k ` is how much up or down is moves/shifts We can see in your problem that ` y = 8(x + 5)^2 ` means that ` a = 8 ` and ` h = -8 ` With how I defined the meanings, you should be able to figure our your options - but if you are still stuck, just let me know